OpenStudy (anonymous):

You stand 40 ft from a tree. The angle of elevation from your eyes (5 ft above the ground) to the top of the tree is 47°. How tall is the tree? Round your answer to the nearest foot.

OpenStudy (anonymous):

height of tree is your height 5ft plus x say x/40 =tan of given angle of 47 deg=1.07 so x=4*1.07=4.29 so tree height add 5 =9.29

OpenStudy (anonymous):

thank you ^.^